Given that Y=asinwt + bt + ct^2coswt. The unit of abc is same as that of{w is omega}a)yb.y/tc.(y/t)^2d.(y/t)^3

Grade:12

1 Answers

Dhruv
6 years ago
unit of a = m       unit of b = m/s      unit of c = m/s2
 
Multiplying a, b and c we get,  unit of abc = m3/s 
 
Hence unit of abc =  unit of  (y/t)3